A plank of mss M is placed on a rough horizontal surface and a constant horizontal force is applied to it. A man of mass m runs on the plank. Find the acceleration of the man so that the plank does not move on the surface. Coefficient of friction between the plank and the surface is µ . Assume sufficient friction between man and the plank to avoid slipping of man on the plank.

Grade:10

1 Answers

AJIT AskiitiansExpert-IITD
13 years ago

Dear Kaushik ,

Two cases can arise :

1) the man runs in the direction of application of the force F on the plank

Then the friction acts on him in the direction of F and so he applies the friction in opposite direction of  F on the plank . His accelecration be a. As he increases a , the force in backward direction increases and eventually the limit reaches when the Friction below on the plank reaches its maximim value = µ (m+M) g

in this case ,  mamax - F =  µ (m+M) g , so

         amax =  µ (m+M) g + F /m

so  a <= (µ (m+M) g + F )/m

2) the man runs opposite to the direction of application of the force F on the plank

Then the friction acts on him in the opposite  direction of F and so he applies the friction in the direction of F on the plank . His accelecration be a. As he increases a , the force in forward  direction increases and eventually the limit reaches when the Friction below on the plank reaches its maximim value = µ (m+M) g.

in this case ,  mamax + F =  µ (m+M) g , so

         amax =  µ (m+M) g - F /m

so  a <= (µ (m+M) g -F )/m

  Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

Askiitians Expert

Ajit Singh Verma IITD